OpenStudy (anonymous):

sin(pi/4+x)=2sqrt(2)(cosx+sinx)

OpenStudy (anonymous):

prove the identity. really need help

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle \sin\left(\frac{\pi}{4}+x\right)=\sin\left(\frac{\pi}{4}\right)\cos(x) +\sin(x)\cos\left(\frac{\pi}{4}+x\right) }\) \(\large\color{black}{ \displaystyle =\sin\left(45^\circ\right)\cos(x) +\sin(x)\cos\left(45^\circ\right) }\) and \(\large\color{black}{ \displaystyle \cos\left(45^\circ\right)=\sin\left(45^\circ\right)=\frac{\sqrt{2}}{2} }\)

OpenStudy (solomonzelman):

go on...

OpenStudy (anonymous):

just the left side?

OpenStudy (anonymous):

sqrt2/2cosx+sinxsqrt2/2

OpenStudy (solomonzelman):

you know the value of cosine and sine of 45 degrees. So use that, and then factor the expression.

OpenStudy (anonymous):

thank you so much I've got it now (:

OpenStudy (solomonzelman):

very good